OpenStudy (anonymous):

Q)A certain quantity of electricity liberates 9g of aluminiun. The volume of oxygen liberated at r.t.p from dilute sulphuric acid by the same quantity of electricity is: A)6 dm^3 B)8 dm^3 C)12 dm^3 D)24 dm^3

OpenStudy (anonymous):

electrolysis = decomposition \[2Al_2O_3\rightarrow4Al+3O_2\] find the no of moles of AL= 1/3 moles 4AL : 3O2 1/3 : x x = 1/4 which is the number of moles of O2 then multiply with 24dm^3 = 6 dm^3

OpenStudy (anonymous):

Where does dilue sulphuric acid come in this? >_<

OpenStudy (anonymous):

i think its 6 whats the answer?

OpenStudy (anonymous):

6 dm^3 is the answer. But why work it out by the decompoition of Al2O3 and not by H2SO4?

OpenStudy (anonymous):

coz even in the decomposition of H2SO4 ionspresent in it = H+/SO4- H+/OH- O2 is dissipated in the same amount, coz the same amount current is used...and u can see the conc of O2 is simmilar to that of aluminum oxide...but if u take it as hydrogen then it would double...

OpenStudy (anonymous):

Okay, thanks!